OpenStudy (anonymous):

suppose you deposit 750$ in a savings account that pays 4.8% interest compounded annually . No withdrawals or additional deposits were done, what id the balance after 6 years ?

OpenStudy (anonymous):

The Compound Interest Equation P = C (1 + r/n) nt where P = future value C = initial deposit r = interest rate (expressed as a fraction: eg. 0.06) n = # of times per year interest in compounded t = number of years invested Source: http://math2.org/math/general/interest.htm

OpenStudy (anonymous):

Sorry, I do this every time. The last part of the equation should be to the power of nt, so P=C(1+r/n)^nt

OpenStudy (anonymous):

It doesn't copy and paste well.

OpenStudy (anonymous):

so whats the balance after 6 years ?

OpenStudy (anonymous):

Simplifies to this when the interest is compounded annually P = C (1 + r)^t

OpenStudy (anonymous):

Now you need to put the information that you have into the equation. You know the initial deposit, the interest rate and the number of years, P = 750(1+0.048)^6

OpenStudy (anonymous):

So, back at you. What do you get for a balance after 6 years?

OpenStudy (anonymous):

@mathdrills of course you are correct. but it says "annually" so the formula is much simpler. just \[P(1+r)^y\]

OpenStudy (anonymous):

lol, I put that formula in after, but usually the next question asks for a different compounding period, so might as well give the standard equation and the website link right away.